### v1

# Highly Undersampled MRI Reconstruction via a Single Posterior Sampling of Diffusion Models

Abstract
Incoherent k-space under-sampling and deep learning-based reconstruction methods have shown great success in accelerating MRI. However, the performance of most previous methods will degrade dramatically under high acceleration factors, e.g., 8 or higher. Recently, denoising diffusion models (DM) have demonstrated promising results in solving this issue; however, one major drawback of the DM methods is the long inference time due to a dramatic number of iterative reverse posterior sampling steps. In this work, a Single Step Diffusion Model-based reconstruction framework, namely SSDM-MRI, is proposed for restoring MRI images from highly undersampled k-space. The proposed method achieves one-step reconstruction by first training a conditional DM and then iteratively distilling this model. Comprehensive experiments were conducted on both publicly available fastMRI images and an in-house multi-echo GRE (QSM) subject. Overall, the results showed that SSDM-MRI outperformed other methods in terms of numerical metrics (PSNR and SSIM), qualitative error maps, image fine details, and latent susceptibility information hidden in MRI phase images. In addition, the reconstruction time for a 320320 brain slice of SSDM-MRI is only 0.45 second, which is only comparable to that of a simple U-net, making it a highly effective solution for MRI reconstruction tasks.
MRI Acceleration, Single-Step Diffusion Model (SSDM), QSM Acceleration, Image Reconstruction
1 Introduction
Magnetic Resonance Imaging (MRI) is a non-invasive and radiation-free imaging modality that has been widely used in clinical diagnosis and research [1, 2] thanks to its excellent tissue contrasts and high image resolutions. However, MRI acquisitions are relatively slow since large areas of k-space need to be traversed, which not only leads to patient discomfort, but also increases the risk of motion artifacts.
Incoherent undersampling combined with compressed sensing (CS) reconstruction algorithms has been substantially studied to accelerate MRI scans, by solving an ill-posed inverse problem. Many traditional iterative reconstruction algorithms have been developed since the pioneering work of Lustig [3, 4], such as image sparsity-based CS methods [3, 4, 5, 6, 7] or low-rankness models [8, 9, 10]. However, the performance of these algorithms will substantially decrease at high acceleration factors, leading to image blurring or residual artifacts. In addition, they are usually computationally expensive and require manual hyperparameter tuning [11].
Deep learning (DL) algorithms have been increasingly popular as an alternative to conventional iterative methods for MRI acceleration. Many end-to-end supervised deep neural networks [12, 13, 14, 15], such as U-net [16] based methods, have been proposed for MRI reconstruction recently, demonstrating better artifact removal and improved fine details, compared with traditional iterative methods. However, the generalization capability of these methods is limited, and their performance will degrade when applied to data out of the training distribution. Whereas, unrolling-based or model-based deep neural network [17, 18] is another framework incorporating physical models in deep neural network designs, demonstrating better results as well as generalization capability. However, these two frameworks will both decrease their performance when applied to highly undersampled k-space data (e.g., 8 or higher).
Recently, diffusion models (DM)[19, 20, 21, 22, 23] or noise-conditioned score networks are increasingly popular for MRI reconstruction at high acceleration factors [24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. This new framework holds great potential for achieving ultra-high k-space acceleration factors, as long as a desirable generative model is successfully trained, which allows high-quality MRI synthesis from only random noise. However, a dramatic number of iterative reverse posterior sampling steps are usually necessary for DM to generate reliable results, which will accumulate reconstruction times, making DM-based MRI reconstruction framework much slower than previous DL methods, e.g., end-to-end network-based methods [12, 13, 14].
In this work, a Single Step Diffusion Model-based method, namely SSDM-MRI, is designed to reconstruct MRI images from highly subsampled k-space in just one posterior sampling step, thus enabling much faster computation than conventional DM-based methods. The proposed method was developed by progressively distilling a pre-trained conditional DM, which was obtained through a supervised learning scheme to enhance the stability and controllability. Comprehensive experiments were conducted on both the public fastMRI brain and knee datasets [34] and an in-house complex-valued brain dataset to compare the proposed SSDM-MRI with multiple established MRI reconstruction methods, including both iterative algorithms and deep learning frameworks. The performances of SSDM-MRI at different acceleration rates were also investigated. In addition, we also evaluated SSDM-MRI’s performance on Quantitative Susceptibility Mapping (QSM) acceleration task using the in-house brain dataset.
2 Theory and Related Works
2.1 Problem Formulation
MRI reconstruction from subsampled k-space data can be formulated as:
where represents the undersampled k-space measurements and is the underlying image, and ; is the measurement noise, and represents the system forward model, with being the undersampling mask and as the Fourier transform. Equation (1) is also known as the general forward model of MRI reconstruction, serving as a training data generator in many supervised learning frameworks [35, 14, 36].
2.2 Conventional Iterative and Deep Learning Algorithms
Equation (1) could be solved using the following optimization framework, which can be calculated iteratively:
where ; is the magnitude images, and is the phase signals; is a loss function enforcing data fidelity, e.g., the commonly used l2 norm; is a manually-designed regularization term, which can be a sparsifying transformation (e.g., Total Variation and Wavelet Transform), and is a weighting hyperparameter. In most previous studies [3, 37, 38, 39], the phase signal is simply estimated from the low-frequency part of k-space to simplify this problem; however, a complex phase regularizer is also necessary for phase-based advanced applications [40, 41, 15], e.g., Quantitative Susceptibility Mapping (QSM).
Deep neural networks, due to their excellent performances in complex feature representation capability compared with manually-designed regularizers (i.e., in Eq. (2)), have demonstrated improved reconstruction performances. End-to-end neural networks [12, 13, 14] and deep unrolling networks [17, 18] are two very commonly used frameworks in CS-MRI. The former can be briefly described as follows:
where is the learnable parameters in the deep network , and is the loss function adopted for network training; represents the techniques used to address the potential overfitting problems in deep and large models. In this scheme, end-to-end networks tried to directly learn the non-linear mapping between training inputs, which is usually the zero-filling reconstructions (i.e., ), and training labels (). Whereas deep unrolling networks, on the other hand, are usually constructed by iteratively stacking physical-model modules and neural networks. The network part usually tries to learn the mapping between latent variables in conventional iterative algorithms (i.e., learning the proximal gradient descent process in ISTA-Net [18]).
2.3 Diffusion Models for MRI Reconstruction
Diffusion models (DM)[19, 21, 22] or the noise-conditioned score networks (NCSN)[20, 23] are new types of generative models that are capable of synthesizing high-quality images via gradually denoising Gaussian noise. It involves a forward noise adding process to convert the original image into pure Gaussian noise, and an iterative backward sampling process to recover the original images in consecutive steps. DM is naturally suitable for solving inverse problems and has shown great potential in MRI reconstructions [24, 25, 26, 27, 28, 29, 30, 31].
Suppose that the desired MRI images follow a specific unknown distribution (i.e., ), then the reconstruction of MRI images () from subsampeld k-space () can be formulated as a posterior sampling process , and the key to obtain reliable samples from this conditional distribution in the DM backward process is to estimate the corresponding score function [20, 42] (i.e., the noise estimation in DM). One can train a deep neural network to estimate this conditional score () in a supervised or unsupervised manner.
2.4 Progressive Diffusion Model Distillation
One major drawback of DM-based methods is the slow sampling (backward) process, which usually takes a large number of posterior sampling steps. Distillation strategies were commonly employed to accelerate the posterior sampling process in many previous works [23, 43], among which progressive distillation [44] has been one of the most effective techniques. It iteratively reduces the number of necessary sampling steps by distilling a teacher diffusion model into a faster student model, which then becomes a new teacher for the next iterative distillation. Every distillation reduces the number of necessary sampling steps from steps to steps.
3 Method
3.1 Overall Framework of SSDM-MRI
The proposed SSDM-MRI takes the zero-filling reconstructions as well as the Gaussian noise as inputs (concatenated along the channel dimension), and generates desired MRI images as outputs, as demonstrated in Fig. 1. To achieve robust image reconstruction with only one posterior sampling, a conditional denoising diffusion probability model (DDPM) [45] was first trained using supervised learning, and it was then progressively distilled to speed up the sampling process, decreasing necessary reverse sampling steps to only one step. Data consistency [46, 15] was also performed to further refine the DM-synthesised images to impose the k-space data fidelity, as described in Eq. (4) below:
where is the original diffusion model reconstructions in k-space, is the undersampling mask, and is the subsampled k-space data. The final reconstruction of SSDM-MRI is the inverse Fourier Transform of k-space signal .
3.2 Diffusion Model Pretraining and Distillation
Similar to most previous works [19, 20], the forward process is a stable Markov chain of diffusion steps, gradually adding Gaussian noise into the fully-sampled MRI images (), which is controlled by a parameter scheduler . In this work, was set as 128 and a cosine noise scheduler () was adopted during DM pretraining, and the forward process can be expressed as follows:
where is the noise contaminated image at time step , is the noise free images, is pure Gaussian noise, and .
The pre-trained conditional DM was obtained by training a self-tailored residual U-net () on a large number of paired datasets () simulated based on Eq. (1), as detailed in Algorithm 1. Overall, the U-net takes both and zero-filling reconstructions () as its major inputs (concatenated along the channel dimension). Similar to most previous U-net models, this adopted U-net consists of an encoding and a decoding path, each consisting of three modules, and each module contains three convolutional blocks with residual connections from the block input to the output. Three pooling layers and three corresponding unpooling layers were included for multiscale or multi-resolution learning. Self-attention operations were only carried out on the latent features of the lowest resolution due to memory constraints. Overall, the network consists of 52 convolutional layers (kernel size: 33), 3 max-pooling layers (kernel size: 22), 3 transposed convolutional layers (kernel size: 22), 51 batch-normalization layers, 77 rectified linear units (ReLUs), 3 concatenations, 25 residual connections, and 6 self-attention operations. The network also takes the noise scheduler () as its third input, which was encoded using a simple thee-layer multi-layer perceptron, inspired by previous works [45] focusing on conditional generation.
In this work, the network is designed to predict a latent variable , which is better suited for distillation [44], considering the fact that the standard noise prediction parameterization used in DDPM is not well suited for distillation, particularly as the signal-to-noise ratio (SNR) approaches zero, where small errors in noise prediction can be significantly amplified, leading to unstable training and degraded performance. Moreover, when distilling down to very few sampling steps, the connection between noise prediction and data reconstruction becomes unreliable. Using the v-diffusion parameterization, the denoising process turns into , which better preserves stable supervision signals during distillation and enables more efficient sampling. Following this modification, progressive distillation is then carried out iteratively to achieve robust image reconstruction to achieve one-step reconstruction. In each distillation iteration, the student model is optimized to halve the overall sampling steps to by learning to replicate the behavior of two sampling steps of the teacher model in just one step, as detailed in Algorithm 2.
3.3 SSDM-MRI for multi-coil data and QSM acceleration
Similar to our previous work [15], the reconstruction of SSDM-MRI for multi-coil data can be carried out as follows: (1) zero-filling reconstruction was first performed for each receiver channel, (2) then, nm the coil sensitivity maps were estimated using the POEM method for coil combination [47], (3) next, the combined images were fed into SSDM-MRI network (without data consistency) for initial reconstruction, (4) then the initial reconstructions were multiplied by the estimated coil sensitivities to convert back into a multi-channel k-space, (5) finally, data consistency was conducted for every coil to obtain images for each coil channel. Finally, these multi-coil images were re-combined using the POEM method [47].
Although the proposed SSDM-MRI is designed for 2D MRI reconstructions, it can be easily transferred to advanced 3D MRI-based applications in a slice-by-slice manner with proper 1D Fourier Transform preprocessing and k-space data consistency, if the acquisition follows the strategy proposed in our previous QSM acceleration work [15]. iQSM+ [48] was adopted for QSM reconstruction from reconstructed MRI phase images.
3.4 Training Data Preparation and Network Training
In this work, two different instances of the proposed SSDM-MRI were obtained by using two different types of training datasets, i.e., (1) the public fastMRI brain dataset [34], and (2) an in-house multi-echo gradient echo (mGRE) complex-valued MRI dataset used in our recent work [15]. In this study, except for the QSM-oriented comparative study, all other evaluations were carried out using the one trained on fastMRI data.
Details about these two datasets are as follows:
(1) fastMRI Brain Dataset We extracted a total of 34,063 2D brain images (size: 320320) from 2,610 volumes randomly selected from the fastMRI brain train dataset batch [34]. The last three slices of each volume were discarded to avoid noise-only slices. All the images were linearly normalized (divided by the maximum intensity of the image) before being fed into the network for training . Different undersampling masks of various types (including 1D and 2D Gaussian masks, and 2D Poisson masks under various AFs, i.e., 4, 6, 8, 10, 12, and 15) were randomly simulated at the beginning of every training mini-batch to synthesize subsampled k-space data and the corresponding zero-filling reconstructions.
(2) mGRE Brain Data A total of 44,063 complex-valued slices (image size: 256256) were retrospectively obtained from 47 subjects (Institutional ethics board approvals have been obtained for all MRI-QSM brain data used in this work), which were acquired at 3T (Discovery 750, GE Healthcare, Milwaukee, WI), with the following parameters: 8 unipolar echoes, first TE / TE = 3 ms / 3.3 ms, TR = 29.8 ms, FOV = 256256128, resolution = 1 mm isotropic. The image normalization and subsampled k-space simulation were made the same as the pipeline described for fastMRI brain dataset.
All network training were conducted on two NVIDIA V100 (32 GB) GPUs, based on the loss functions defined in Algorithm 1 and 2. For DM pretraining, all learnable parameters in the network were initialized as random numbers with a mean of 1 and standard deviation of 0.001, and are optimized using the Adam optimizer with a learning rate of 5 for 200 epochs. The proposed models were implemented using Pytorch 1.13.1. The source codes and the trained models are available at https://github.com/YangGaoUQ/SSDM-MRI.
4 Experiments
4.1 Evaluation Datasets
The performance and generalization capability of the proposed SSDM-MRI were comprehensively evaluated on the following retrospectively and prospectively undersampled data, including:
128 in distribution single-coil brain slices from fastMRI at acceleration factors (AF) of 8 and 12 (1D Gaussian subsampling, ACS = 12 and 6 lines, respectively).
Out-of-distribution (OOD) fastMRI knee dataset (156 single-coil images) at AF of 12 (Gaussian 2D).
A single-coil pathological slice containing brain lesions (AF = ), simulated from a subject acquired at 3T using the parameters of the mGRE brain training dataset.
Retrospective multi-coil brain (78 images at AF=12) and multi-coil knee (103 images at AF=15) datasets with ground truths.
Two 12-coil prospectively undersampled brain and knee slices (AF = and ) randomly selected from fastMRI testing dataset.
A brain data (size: ) retrospectively undersampled at AF = 8 using sampling masks adopted in [15] to evaluate SSDM-MRI’s performance on advanced QSM acceleration task.
4.2 Performance Evaluation and Numerical Metrics
P-LORAKS | CCPC | ADMM-Net | U-net | Score-MRI | SSDM-MRI |
2.67 s | 221.45 s | 1.91 s | 0.36 s | 315.18 s | 0.45 s |
P-LORAKS and CSPC were run on a Intel i5-12400 CPU Processor, while the other deep learning method were evaluated on a NVIDIA 3060 GPU.
SSDM-MRI was compared with the two traditional algorithms and three deep learning methods: (1) scan-specific P-LORAKS [49], (2) Compressed Sensing with Phase Cycling (CSPC) [40], (3) the unrolling-based ADMM-Net [17], (4) end-to-end U-net based method [16], and (5) DM-based Score-based MRI [26, 29] ( overall sampling steps) on the above described datasets of different acceleration factors and subsampling patterns. The deep learning methods were re-trained using the same datasets that were originally used for training the SSDM-MRI. The reconstruction times of different methods on a image of size are summarized in Table 1, indicating that the reconstruction speed of SSDM-MRI is much faster than that of Score-MRI, since only one reverse step is needed.
Performance evaluation experiments are carried out as follows: (1) The proposed SSDM-MRI was first compared with other reconstruction methods using the single-coil fastMRI brain dataset. (2) Then, it was tested on OOD fastMRI knee images and the simulated pathological lesion brain image to test its generalization capability. (3) Next, SSDM-MRI’s performances were further validated on both the retrospective and prospective multi-coil fastMRI data. (4) In addition, SSDM-MRI’s performance was evaluated on one 8-echo GRE brain for QSM acceleration following [15], which could compare various methods at the whole brain level instead of a single 2D slice, and evaluate their capability of preserving latent susceptibility information. (5) One more simulated study was carried out to investigate the performance of SSDM-MRI against various acceleration factors (AF = 4, 6, 8, 10, 12, respectively). (6) Finally, the stability of SSDM-MRI was compared with Score-MRI by inspecting the generation of uncertainty maps.
P-LORAKS | ADMM-NET | U-net | Score-MRI | SSDM-MRI | |||
Single-coil knee (156 images) | Gaussian 2D | 25.94 ± 2.031 | 25.56 ± 2.074 | 27.94 ± 2.343 | 27.46 ± 2.265 | 29.95 ± 2.144 | 29.49± 2.213 |
×8 | 0.76 ± 0.047 | 0.76 ± 0.044 | 0.79 ± 0.043 | 0.79± 0.056 | 0.81 ± 0.071 | 0.82 ± 0.067 | |
Gaussian 1D | 21.74 ± 1.754 | 21.27 ± 1.337 | 27.56 ± 2.318 | 25.96 ± 2.866 | 28.19 ± 2.879 | 28.08 ± 2.311 | |
×12 | 0.56 ± 0.081 | 0.54 ± 0.076 | 0.77 ± 0.023 | 0.75 ± 0.085 | 0.76 ± 0.023 | 0.78± 0.043 | |
Single-coil brain (128 images) | Gaussian 1D | 25.34 ± 1.613 | 25.11 ± 1.431 | 32.93 ± 2.557 | 31.71 ± 2.808 | 32.77 ± 1.995 | 33.83 ± 2.311 |
×8 | 0.82 ± 0.049 | 0.82 ± 0.062 | 0.92 ± 0.017 | 0.91 ± 0.038 | 0.92 ± 0.024 | 0.93± 0.029 | |
Poisson 2D | 24.77± 1.983 | 24.54 ± 2.454 | 29.11 ± 2.353 | 28.98 ± 1.967 | 29.29 ± 2.348 | 30.19 ± 2.103 | |
×15 | 0.70 ± 0.061 | 0.69 ± 0.051 | 0.86 ± 0.037 | 0.84 ± 0.066 | 0.88 ± 0.038 | 0.89 ± 0.034 | |
Multi-coil brain (78 images) | Gaussian 1D | 21.45 ± 1.217 | 20.71 ± 1.146 | 28.15 ± 2.254 | 27.71 ± 2.365 | 29.59 ± 1.897 | 31.28 ± 2.314 |
×12 | 0.63 ± 0.084 | 0.63 ± 0.078 | 0.85 ± 0.024 | 0.83 ± 0.028 | 0.89 ± 0.047 | 0.90 ± 0.031 | |
Random 1D | 20.41 ± 1.943 | 19.85 ± 2.575 | 26.14 ± 2.154 | 25.42 ± 2.993 | 26.67 ± 2.079 | 27.66 ± 2.473 | |
×12 | 0.54 ± 0.074 | 0.53 ± 0.066 | 0.78 ± 0.061 | 0.77 ± 0.083 | 0.85 ± 0.053 | 0.86 ±0.041 | |
Multi-coil knee (103 images) | Poisson 2D | 23.67± 2.185 | 22.57 ± 2.553 | 28.31 ± 2.438 | 27.98 ± 2.567 | 29.35 ± 2.379 | 29.89 ± 2.132 |
×15 | 0.62 ± 0.072 | 0.63 ± 0.063 | 0.75 ± 0.087 | 0.74 ± 0.076 | 0.77 ± 0.073 | 0.78 ± 0.067 | |
Gaussian 1D | 20.31 ± 2.743 | 20.15 ± 2.677 | 26.71 ± 2.791 | 26.42 ± 2.990 | 27.97 ± 2.673 | 29.01 ± 2.510 | |
×12 | 0.55 ± 0.078 | 0.54 ± 0.076 | 0.73 ± 0.074 | 0.71 ± 0.083 | 0.78 ± 0.053 | 0.78 ± 0.057 |
Peak Signal to Noise Ratio (PSNR) and Structural Similarity (SSIM) [50] were reported for quantitative assessments of different methods on magnitude images, while PSNR and XSIM (a QSM-specific SSIM variants) [51] were reported for the QSM acceleration task. In addition, voxel-wise linear regressions on QSM values of Substantia Nigra (SN) and Red Nucleus (RN) regions were carried out to compare different methods.
5 Results
5.1 Evaluation on single-coil retrospective data
Fig. 2 compares SSDM-MRI with various MRI accelerating methods on two in-distribution fastMRI single-coil brain slices (AF = 12 and 8, respectively). Overall, SSDM-MRI showed the highest PSNR and SSIM on both datasets, and the minimum errors in the zoomed-in region. P-LORAKS and CSPC showed substantial residual artifacts (red arrows), which are absent in Score-MRI and SSDM-MRI. Meanwhile, SSDM-MRI preserved finer structures (green arrows) compared with other methods.
SSDM-MRI’s performance is evaluated on two OOD images (i.e., a fastMRI knee and a pathological brain) in Fig. 3. SSDM-MRI resulted in the minimum error maps and the best fine image details for the knee slice, as pointed out by the green arrow. For the pathological brain image, SSDM-MRI led to the best metrics among all methods, while ADMM-Net led to the minimum error map on the brain lesion.
5.2 Evaluation on multi-coil fastMRI data
5.2.1 Evaluation on retrospective data
The proposed SSDM-MRI is compared with P-LORAKS, CSPC, ADMM-Net, U-net, and Score-MRI on two retrospective fastMRI brain and knee images (12-coil) were compared in Fig. 4 at AF = 12 and 15, respectively. P-LORAKS and CSPC led to apparent artifacts and reconstruction errors at these high AFs, while U-net resulted in over-smoothed results. SSDM-MRI not only led to the best numerical metrics but also showed the best fine structures (green arrows), compared with other methods.
Table 2 summarizes the performances of different methods on four retrospective dataset obtained from fastMRI: (1) single-coil brain, (2) single-coil knee, (3) multi-coil brain, and (4) multi-coil knee images at different AFs and different types of sampling masks. SSDM-MRI achieved on average the best PSNR and SSIMs in nearly all situations, compared with the other algorithms.
5.2.2 Evaluation on prospective data
Different methods for two prospectively undersampled data are compared in Fig. 5. Similar to the simulation case, traditional algorithms, ie., P-LORAKS, CSPC demonstrated substantial artifacts (red arrows), which were successfully removed by ADMM-Net, U-net, Score-MRI, and SSDM-MRI. In addition, SSDM-MRI also significantly reduced the image blurring apparent in U-net and demonstrated the best fine structures as pointed out by the green arrows.
5.3 Comparison on QSM acceleration task
Fig. 6 compares different methods on a QSM brain volume at 8 acceleration. While all methods showed dramatic detail loss, compared to the fully-sampled QSM ground truth. SSDM-MRI resulted in the minimum error map and the best PSNR and XSIM (38.80 and 0.66). Linear regression was also performed on susceptibility values of various reconstruction methods against those of the ground truth, as shown Fig. 6(b). It confirmed that the proposed SSDM-MRI achieved the most accurate quantitative measurements in both SN ( =0.93) and RN ( =0.81) regions.
5.4 The effects of accelerating factors
The reconstructions of SSDM-MRI under 5 different AFs are shown in Fig. 7, on one retrospectively subsampled brain and a knee image. None of the reconstructed images showed apparent artifacts; however, as the AFs increases, more fine details gradually became blurred, which was also confirmed in the error maps.
5.5 Reconstruction stability of SSDM-MRI
Fig. 8 illustrates the uncertainty maps of SSDM-MRI and Score-MRI of 5 repetitive reconstructions on one knee slice at AF = 4 , 8, 12 respectively. As the AF increases from 4 to 12, both DM-based methods exhibited a notable increase in uncertainty images, with the Score-MRI demonstrating slightly larger standard deviations than the proposed SSDM-MRI.
6 Discussion and Conclusion
In this work, we demonstrated the feasibility of MRI reconstructions from highly undersampled k-space with only a single-step posterior sampling of a diffusion model (DM). Specifically, an SSDM-MRI framework was constructed by progressively distilling a conventional DM. Extensive comparative studies were carried out to compare the performances of SSDM-MRI with those of multiple existing algorithms, i.e., P-LORAKS [49], CSPC [40], ADMM-Net [17], U-net [16], and DM-based Score-MRI [26] on both public fastMRI images and a private QSM subject. In general, the proposed method exhibited the best results on most images with fewer errors, improved numerical metrics and image fine details, and more accurate susceptibility measurements in the substantia nigra and red nucleus.
There have already been several DM-based methods proposed for MRI acceleration [24, 25, 26, 27, 28, 29, 30, 31], showing better reconstruction performances, especially when compared with traditional iterative and existing end-to-end deep learning (DL) based methods [12, 13, 14]. However, the computational costs of DM-based methods are usually much higher than previous DL methods, e.g., U-net-based frameworks, due to the need of a relatively large number of poster sampling steps, making the overall reconstruction time much longer and hindering their applications in clinical practice. To the best of our knowledge, SSDM-MRI is the first single-step DM-based method, whose reconstruction time is comparable to the end-to-end deep neural networks, as evident in Table 1.
In this work, the proposed SSDM-MRI demonstrated promising generalization capability on the OOD knee data, despite the fact that it was trained on brain images only, demonstrating SSDM-MRI’s ability to learn the conditional prior underlaid in shared anatomical and statistical patterns of brain and knee images instead of simply memorizing the features. Although the adaptability of SSDM-MRI may not be as good as fully unsupervised methods (e.g., the slight performance decrease in Tabel 2 for knee images), it still demonstrates great robustness across vairous anatomies.
In the QSM experiment, different methods were repetitively tested on 1024 (128 slice number 8 echo number) consecutive 2D images, covering the whole brain before QSM reconstructions from MRI phases. The proposed SSDM-MRI not only achieved on average the best results for all brain slices, but also preserved the best slice-to-slice coherence. It also resulted in the best QSM measurements, demonstrating its capability of preserving latent and weak susceptibility signal buried in the MRI phase images.
SSDM-MRI’s performances were also evaluated against varying accelerating rates, and the results showed that no substantial artifacts were introduced. However, the image quality also gradually degraded with increasing AFs, and in the future, we will continue working on obtaining high-quality images with high AFs. In addition, the reconstruction stability also needs to be further optimized to obtain more reliable solutions.
In conclusion, this study proposed a single-step diffusion model(DM)-based method, SSDM-MRI, for MR image reconstruction from highly undersampled k-space data, via a one-step posterior sampling from a conditional DM. Experimental results demonstrated that SSDM-MRI consistently led to improved results compared with several state-of-the-art methods on both magnitude and QSM reconstruction tasks, respectively, in a much shorter reconstruction time.